# Inference-time Scaling of Diffusion Models through Classical Search

### v1

@@ 6 Limitations and Conclusion
Limitations
Despite the performance improvement of inference-time scaling, it still requires more hyperparameter tuning than training-based methods. One approach is to design an adaptive hyperparameter tuning framework, which could increase or decrease automatically based on the current sample. For example we can tune the threshold of DFS automatically given the pass rate of samples. Another potential concern is that the generated samples may be used as adversarial examples to attack verifiers. Our method can be used to search for weaknesses of verifiers, which in turn can be used to train a more robust verifier.
Conclusion
In this work we propose a principled framework for inference-time scaling through classic search principles, unlocking the potential of scaling inference-time compute of diffusion models in a variety of downstream tasks. Inference-scaling can also be applied to training models via expert iteration, generating high-quality synthetic data beyond the model’s ability. We leave this as a interesting future direction.
@@ Appendix B Additional Related Works
Inference-time scaling
@@
Scaling compute in inference-time with “slow thinking” has its long history grounded in cognitive science, known as “system 2” thinking [32, 63]. In [52, 53], Newell and his colleagues formalized problem solving as tree search in a combinatorial space, and [93] uses tree-of-thoughts to enable LLM reasoning with multiple exploration paths, using BFS and DFS as strategic search algorithms.
Recently, long chain-of-though (CoT) reasoning has demonstrated remarkable performance for LLM reasoning [30, 21], where the long CoT reasoning ability is incentivized through reinforcement learning [21]. Notably, the CoT process demonstrates reasoning activities such as self-verification, backtracking and self-correction. Using a process reward model [39], we can also conduct explicit tree search without training the language model. [85] propose reward-balanced search (REBASE) which is a special instance of BFS, and [64] applied beam-search to difficult math problems, showing compute-optimal inference can be achieved via selecting different strategy for problems with different difficulty. We refer the readers to [79] for a comprehensive review.
Inference-time scaling could also be used to improve the model itself, known as expert iteration. In [32], they proposed reinforcement learning with expert iteration in games, where the expert is constructed combining the base policy with Monte Carlo Tree Search (MCTS), and [86] applied expert iteration to automated theorem proving. Self improvements can also be achieved through iterative self-training [69, 96, 13, 61].
Inference scaling in diffusion models
The inference-time compute of diffusion models depends heavily on the number of denoising steps. [91] showed that recursive restart sampling can reduce cumulative error during sampling, which can be regarded as scaling the number of denoising steps. More recently, [62] proposed a Sequential Monte Carlo (SMC) [84, 99] style method, known as Feynman-Kac steering, which can be seen as a instance of BFS. Besides image generation, they also applied it to diffusion language models [23, 59]. Additionally, [48] explored inference-time scaling of diffusion models with local zero order search and global search over paths for image generation. They also experimented with different verifiers, such as oracle verifiers, self-supervised verifiers, and studied the verifier-task alignment problem. There zero-order search can be understood as an uninformed version of local search which exhibits low efficiency, and when utilizing gradients, they need to back-propagate through the entire diffusion sampling chain, causing high computation and memory overheads. Compared with their work, we propose efficient gradient-based local search with theoretically grounded Langevin MCMC, which we show is crucial in many tasks. Also, we provide systematic experiments on the compute efficiency of global search methods. Their proposed methods can thus be understood as a instance within our search framework.
Apart from search, [38] exploits the in-context learning abilities of foundation models to provide revision during sampling. Specifically, they leverage the multi-modal capabilities of VLMs to provide feedbacks on past generated images, and train the model to condition on past images and feedback.
Diffusion models and applications
Diffusion models [26, 68, 76] has shown great performance in generative modeling for continuous data domains, such as image [18, 12], videos [28] and molecules [90]. Due to its expressive power on modeling multimodal and complex distributions, they have also been widely used as a decision prior in robotics. [31] proposes the first work on using diffusion models to generate plans. [10] uses the diffusion model for visuo-motor policy in robotics. Recently, a series of robotics foundation models utilize diffusion heads as action experts [5, 73], while [43] trains a end-to-end diffusion transformer for bimanual manipulation. In this work, we demonstrate that inference-time scaling can be especially helpful for decision making tasks with diffusion models.
Control and alignment of diffusion models
To align the diffusion model with flexible objectives, training-free guidance [94, 4, 11, 66, 95, 25] and compositional generation [14, 42] combines the diffusion models with classifiers or other diffusion models at inference time, while RL-based methods [98, 78, 89, 6] finetune the diffusion model using reward or preference signals. ControlNet [97] style approaches have also been used to add additional conditions for sampling, where [71] designs a control block for diffusion transformers, and [88] uses a combination of guidance and controlnet for controllable human motion generation.
@@ C.1 Langevin MCMC as gradient flow in measure space
Following [83], the Langevin SDE in sample space corresponds to gradient flow of the KL-divergence in measure space. Here we provide a brief overview.
@@
Previous works [17, 9] show that ULA can efficiently converge to the target measure $\nu$ if $\nu$ is log-concave and smooth. However, when facing complex and multimodal distributions, we can only guarantee convergence to the concave vicinity.
@@ C.2 Annealed Langevin MCMC Sampling
Langevin MCMC have been used to perform implicit sampling in energy-based models [16] and score-based models [67]. However, these methods suffer from inaccurate score estimation and low density regions [67]. In [67] they propose to perturb the data with gaussian noise, eventually smoothing the data distribution:
@@
and creating a sequence of annealed distributions $\left\{q({\bm{x}}_{t})\right\}_{t=0}^{T}$ which converges to $p_{0}({\bm{x}}_{0})$. Since they are smoothed by gaussian noise, we can improve the mixing time of Langevin MCMC on multimodal distributions by sampling from these intermediate distributions, sharing similar spirits with simulated annealing [33].
In [14], they extend this method to compositional generation of diffusion models. Specifically, we consider sampling from a product distribution $p^{\text{prod}}_{0}({\bm{x}}_{0})\propto p_{0}^{1}({\bm{x}}_{0})p_{0}^{2}({\bm{x}}_{0})$, where $p_{0}^{1}({\bm{x}}_{0})$ and $p_{0}^{2}({\bm{x}}_{0})$ are distributions of different diffusion models. Since we have access to the score functions $\nabla_{{\bm{x}}_{t}}\log q_{t}^{1}({\bm{x}}_{t})$ and $\nabla_{{\bm{x}}_{t}}\log q_{t}^{2}({\bm{x}}_{t})$ through the diffusion model, we can construct a sequence of annealing distributions $\tilde{q}_{t}^{\text{prod}}({\bm{x}}_{t})$ such that:
@@ C.3 Annealed Langevin MCMC with recurrent training-free guidance
In this section, we prove the connection between annealed Langevin MCMC (Alg. 1) and training-free guidance (Alg. 2) in Proposition. 1. We divide the proof into two parts. In Sec. C.3.1 we prove the equivalence between naive recurrence steps and Langevin MCMC. Then in Sec. C.3.2, we prove that adding the guidance term is defining an annealing path that biases towards high score regions.
@@ C.3.1 Equivalence between Langevin MCMC and naive recurrence
Consider the diffusion process with the following stochastic interpolant [47]:
@@
When using DDIM [65] sampler, we have the reverse step as:
@@
where $1+e^{\lambda_{t}-\lambda_{t-1}}\rightarrow 2$ when $T\rightarrow\infty$ and denoising step size approaches 00, as $\lambda_{t}-\lambda_{t-1}\rightarrow 0$.
@@
In DDPM [26], we parametrize the posterior distribution as:
@@
[26] parameterizes the posterior variance as $\Sigma_{\theta}({\bm{x}}_{t},t)=\beta_{t}\bm{I}$ or $\Sigma_{\theta}({\bm{x}}_{t},t)=\tilde{\beta}_{t}\bm{I}$:
@@
while [54] propose to train the posterior variance as $\Sigma_{\theta}({\bm{x}}_{t},t)=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right)$.
@@
where $\Sigma_{\theta}({\bm{x}}_{t},t)\rightarrow\beta_{t}\bm{I}$ when $T\rightarrow\infty$, and the denoising step size approaches 00.
@@
${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+a_{t}s({\bm{x}}_{t}^{i},t)+b_{t}{\bm{\epsilon}}^{i}\,.$ (12)
with $a_{t}\rightarrow 0$, $b_{t}\rightarrow 0$ and $2a_{t}\rightarrow b_{t}^{2}$ as the denoising step size approaches 00. Thus, it can be seen as a approximation of the ULA in Eq. 8, and also a discretization of the Langevin SDE in Eq. 7.
@@ C.3.2 Annealed Langevin MCMC with guidance
When applying training free guidance [94] during the recurrence, we have:
${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+a_{t}s({\bm{x}}_{t}^{i},t)+b_{t}{\bm{\epsilon}}^{i}+\bm{\Delta}({\bm{x}}_{t},t)\,,$
@@
where the second Equation follows from Lemma 3.3 in [94]. Thus, there exists a set of functions $\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})$ such that $\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})=\frac{\sigma_{t}^{2}}{\bm{\Sigma}_{0|t}}\nabla_{{\bm{x}}_{t}}\log f({\bm{x}}_{0|t})$, and we can see that when $t\rightarrow 0$,$\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})=\nabla_{{\bm{x}}_{0}}\log f({\bm{x}}_{0})$. If we additionally incorporate the ‘implicit dynamics’ in Line. 4, our arguments still stands since the smoothed objective $\tilde{f}({\bm{x}})=\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}f({\bm{x}}+\bar{\gamma}{\sigma_{t}}\bm{\delta})$ converges to $f$ with $t\rightarrow 0$ and $\sigma_{t}\rightarrow 0$.
@@
${\bm{x}}_{t}^{i+1}$ $={\bm{x}}_{t}^{i}+a_{t}s({\bm{x}}_{t}^{i},t)+b_{t}{\bm{\epsilon}}^{i}+c_{t}\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}({\bm{x}}_{t})\,,$
which, in the continuous limit of $a_{t},b_{t},c_{t}\rightarrow 0$, converges to the Langevin SDE:
$d{\bm{x}}_{t}^{i}=\nabla_{{\bm{x}}_{t}^{i}}\log\left(q_{t}({\bm{x}}_{t}^{i})\hat{f}_{t}({\bm{x}}_{t}^{i})\right)di+\sqrt{2}d\bm{w}^{i}\,,$
where $i$ is the time index of recurrence steps. Thus, we have defined the annealing path as $\tilde{q}_{t}({\bm{x}}_{t})=q_{t}({\bm{x}}_{t})\hat{f}_{t}({\bm{x}}_{t})\,,~{}t=1,2,\cdots,T$.

### v2

@@ 6 Limitations and Conclusion
In this work, we present a unified and principled framework for inference-time scaling of diffusion models, jointly scaling local and global search. A potential limitation of our approach is that inference-time scaling still requires additional hyperparameters. To address this, heuristics such as evolutionary search can be adopted during hyperparameter tuning, which we leave for future work.
@@ Appendix B Additional Related Works
B.1 Discussions on Concurrent Works
Adaptive Bi-directional Cyclic Diffusion (ABCD) (Lee et al., 2025) introduces a search-based inference-scaling framework that combines elements of both DFS and BFS. Unlike DFS, which relies on a quality threshold to determine backtracking, ABCD maintains a set of particles and backtracks by propagating them across all noise levels. The termination criterion is based on whether backtracking to higher noise levels improves sample quality. Compared with DFS, ABCD achieves smaller score estimation errors since particles are evaluated after full denoising, and it explores the generative space more thoroughly by combining BFS and DFS. However, ABCD cannot adaptively allocate compute across different instances due to its fixed termination condition, and it expends unnecessary compute on easier instances because it denoises a full set of particles regardless of sample quality.
EvoSearch (He et al., 2025) proposes evolutionary search to scale inference compute for image and video generation. At selected timesteps, particles are fully denoised; high-scoring particles are retained, while low-scoring ones are perturbed via noise injection. This method improves over the FK-steering baseline (Singhal et al., 2025a).
In contrast, our work develops a uniform and principled framework grounded in classical search heuristics. BFS and DFS serve as fundamental building blocks for more advanced search methods, and, crucially, we are the first to scale local search using annealed Langevin MCMC. This local search can be seamlessly combined with global search strategies, a joint scaling approach that we show is critical for success in challenging decision-making domains.
B.2 Inference-Time Scaling with Classical Search
We now review the use of classical search for inference-time scaling, including applications beyond diffusion models.
@@ B.2 Inference-Time Scaling with Classical Search
Inference-time compute scaling, often viewed as “slow thinking,” has long-standing roots in cognitive science, where it corresponds to “system 2” reasoning (Kahneman, 2011; Sloman, 1996). In early work, Newell et al. (1959; 1972) formalized problem solving as tree search in combinatorial spaces. More recently, Yao et al. (2023) proposed tree-of-thoughts reasoning for LLMs, enabling exploration along multiple reasoning paths using BFS and DFS as strategic search primitives.
B.3 Inference Scaling in Diffusion Models without Differentiable Optimization
We next provide a more detailed overview of gradient-free inference-scaling approaches in diffusion models.
Inference-time compute for diffusion models is largely determined by the number of denoising steps. Xu et al. (2023b) proposed recursive restart sampling to mitigate cumulative errors, effectively scaling the number of denoising steps. More recently, Singhal et al. (2025a); Kim et al. (2025) introduced Sequential Monte Carlo (SMC) approaches (Wu et al., 2023; Zhao et al., 2024), interpretable as BFS-style algorithms. Their implementations differ mainly in the choice of scoring functions and resampling strategies, with Kim et al. (2025) proposing an increasing tempering schedule. Ma et al. (2025) further explored inference-time scaling for image generation via a combination of local zero-order search and global path search, experimenting with different verifiers (oracle, self-supervised, and task-aligned) and analyzing verifier-task alignment. However, they did not systematically evaluate and improve the compute efficiency of search methods, while we demonstrate significantly improved efficiency with our novel design in search methods. Additionally, Yoon et al. (2025) proposed a Monte Carlo Tree Search-based diffusion framework tailored for trajectory synthesis, which denoises different trajectory segments at distinct steps.
Beyond search-based methods, Li et al. (2025) leverage the in-context learning abilities of foundation models to enable revision during sampling. Their approach exploits multimodal VLM capabilities to provide feedback on intermediate generations and trains the model to condition on both prior samples and corresponding feedback.
B.4 Gradient Guidance of Diffusion Models
We review guidance methods for diffusion models that leverage gradients from verifiers or classifiers.
To sample from a conditional distribution
$\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})\,p(c|{\bm{x}}_{0}),$
where $p$ denotes a classifier, one can define a conditional diffusion process $\{\tilde{q}_{t}({\bm{x}}_{t})\}_{t=0}^{T}$ with initial distribution $\tilde{p}_{0}$. The corresponding conditional score function is
$\nabla_{{\bm{x}}_{t}}\log\tilde{q}_{t}({\bm{x}}_{t})=\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})+\nabla_{{\bm{x}}_{t}}\log p_{t}(c|{\bm{x}}_{t}),$
where $p_{t}(c|{\bm{x}}_{t})=\mathbb{E}_{{\bm{x}}_{0}|{\bm{x}}_{t}}[p(c|{\bm{x}}_{0})]$. Classifier guidance (Dhariwal & Nichol, 2021) trains a noise-dependent classifier $p_{t}(c|{\bm{x}}_{t})$ on noisy samples, using data from the diffusion model together with the base classifier. However, this training procedure is computationally expensive and impractical when the base diffusion model cannot generate target-class samples.
Training-free guidance methods (Ye et al., 2024; Chung & Ye, 2022; Song et al., 2023; Yu et al., 2023; He et al., 2023) approximate the intractable $p_{t}(c|{\bm{x}}_{t})$ with $p(c|{\bm{x}}_{0|t})=p(c|\mathbb{E}_{{\bm{x}}_{0}|{\bm{x}}_{t}}[{\bm{x}}_{0}])$, which can be interpreted as a first-order approximation. These methods introduce bias, though recurrence has been shown to mitigate out-of-distribution artifacts and improve performance in classifier-guided generation (Ye et al., 2024). Nevertheless, the theoretical foundations and scaling behavior of recurrent guidance remain largely unexplored.
In this work, we instead adopt Langevin MCMC sampling in place of score-based ODE sampling. This approach guarantees asymptotic exactness using only the base verifier, without requiring additional training. Relatedly, Du et al. (2024) proposed annealed energy-based models (EBMs) combined with iterative MCMC for inference scaling and reasoning, but their method cannot be applied to pretrained diffusion models and verifiers. For the first time, we establish a connection between annealed Langevin MCMC and recurrent guidance, thereby enabling inference scaling with pretrained foundation models.
@@ C.1 Langevin MCMC as gradient flow in measure space
Following Wibisono (2018), the Langevin SDE in sample space corresponds to gradient flow of the KL-divergence in measure space. Here we provide a brief overview.
@@
Previous works (Durmus & Moulines, 2015; Cheng & Bartlett, 2018) show that ULA can efficiently converge to the target measure $\nu$ if $\nu$ is log-concave and smooth. However, when facing complex and multimodal distributions, we can only guarantee convergence to the concave vicinity.
@@ C.2 Annealed Langevin MCMC Sampling
Langevin MCMC have been used to perform implicit sampling in energy-based models (Du & Mordatch, 2019) and score-based models (Song & Ermon, 2020). However, these methods suffer from inaccurate score estimation and low density regions (Song & Ermon, 2020). In Song & Ermon (2020) they propose to perturb the data with gaussian noise, eventually smoothing the data distribution:
@@
and creating a sequence of annealed distributions $\left\{q({\bm{x}}_{t})\right\}_{t=0}^{T}$ which converges to $p_{0}({\bm{x}}_{0})$. Since they are smoothed by gaussian noise, we can improve the mixing time of Langevin MCMC on multimodal distributions by sampling from these intermediate distributions, sharing similar spirits with simulated annealing (Kirkpatrick et al., 1983).
In Du et al. (2023), they extend this method to compositional generation of diffusion models. Specifically, we consider sampling from a product distribution $p^{\text{prod}}_{0}({\bm{x}}_{0})\propto p_{0}^{1}({\bm{x}}_{0})p_{0}^{2}({\bm{x}}_{0})$, where $p_{0}^{1}({\bm{x}}_{0})$ and $p_{0}^{2}({\bm{x}}_{0})$ are distributions of different diffusion models. Since we have access to the score functions $\nabla_{{\bm{x}}_{t}}\log q_{t}^{1}({\bm{x}}_{t})$ and $\nabla_{{\bm{x}}_{t}}\log q_{t}^{2}({\bm{x}}_{t})$ through the diffusion model, we can construct a sequence of annealing distributions $\tilde{q}_{t}^{\text{prod}}({\bm{x}}_{t})$ such that:
@@ C.3 Annealed Langevin MCMC with recurrent training-free guidance
In this section, we prove the connection between annealed Langevin MCMC (Alg. 1) and training-free guidance (Alg. 2) in Proposition. 1. We divide the proof into two parts. In Sec. C.3.1 we prove the equivalence between naive recurrence steps and Langevin MCMC. Then in Sec. C.3.2, we prove that adding the guidance term is defining an annealing path that biases towards high score regions. Finally, we provide a rigorous convergence analysis in Sec. C.3.3.
@@ C.3.1 Equivalence between Langevin MCMC and naive recurrence
Consider the diffusion process with the following stochastic interpolant (Ma et al., 2024):
@@
When using DDIM (Song et al., 2020a) sampler, we have the reverse step as:
@@
where $1+e^{\lambda_{t}-\lambda_{t-1}}\rightarrow 2$ when $T\rightarrow\infty$ and denoising step size approaches 0, as $\lambda_{t}-\lambda_{t-1}\rightarrow 0$.
@@
In DDPM (Ho et al., 2020), we parametrize the posterior distribution as:
@@
Ho et al. (2020) parameterizes the posterior variance as $\Sigma_{\theta}({\bm{x}}_{t},t)=\beta_{t}\bm{I}$ or $\Sigma_{\theta}({\bm{x}}_{t},t)=\tilde{\beta}_{t}\bm{I}$:
@@
while Nichol & Dhariwal (2021) propose to train the posterior variance as $\Sigma_{\theta}({\bm{x}}_{t},t)=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right)$.
@@
where $\Sigma_{\theta}({\bm{x}}_{t},t)\rightarrow\beta_{t}\bm{I}$ when $T\rightarrow\infty$, and the denoising step size approaches 0.
@@
${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+a_{t}r_{t}s({\bm{x}}_{t}^{i},t)+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}\,.$ (13)
with $a_{t}\rightarrow 0$ and $r_{t}\rightarrow 1$ as the denoising step size approaches 0:
  • For DDIM sampler, we have $a_{t}=\frac{1}{2}\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)$ and $r_{t}=\frac{2}{1+e^{\lambda_{t}-\lambda_{t-1}}}$.
  • For DDPM sampler, we have $a_{t}=\frac{1}{2}\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)$ and $1\leq r_{t}\leq\frac{2}{1+\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}}$.
Thus, it can be seen as a approximation of the ULA in Eq. 9, and also a discretization of the Langevin SDE in Eq. 8.
@@ C.3.2 Annealed Langevin MCMC with guidance
When applying training free guidance (Ye et al., 2024) during the recurrence, we have:
${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+a_{t}r_{t}s({\bm{x}}_{t}^{i},t)+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}+\bm{\Delta}({\bm{x}}_{t},t)\,$,
@@
where the second Equation follows from Lemma 3.3 in Ye et al. (2024). Thus, there exists a set of functions $\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})$ such that $\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})=\frac{\sigma_{t}^{2}}{\bm{\Sigma}_{0|t}}\nabla_{{\bm{x}}_{t}}\log f({\bm{x}}_{0|t})$, and we can see that when $t\rightarrow 0$$\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})=\nabla_{{\bm{x}}_{0}}\log f({\bm{x}}_{0})$. If we additionally incorporate the ‘implicit dynamics’ in Line. 4, our arguments still stands since the smoothed objective $\tilde{f}({\bm{x}})=\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}f({\bm{x}}+\bar{\gamma}{\sigma_{t}}\bm{\delta})$ converges to $f$ with $t\rightarrow 0$ and $\sigma_{t}\rightarrow 0$.
@@
${\bm{x}}_{t}^{i+1}$ $={\bm{x}}_{t}^{i}+a_{t}r_{t}s({\bm{x}}_{t}^{i},t)+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}+c_{t}\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}({\bm{x}}_{t})$
$={\bm{x}}_{t}^{i}+a_{t}r_{t}\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})\hat{f}_{t}({\bm{x}}_{t})^{c_{t}/a_{t}r_{t}}+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}\,$,
Thus, we have defined the annealing path as $\tilde{q}_{t}({\bm{x}}_{t})=q_{t}({\bm{x}}_{t})\hat{f}_{t}({\bm{x}}_{t})^{c_{t}/a_{t}r_{t}}\,,~t=1,2,\cdots,T$.
C.3.3 Convergence analysis
In this section, we provide a rigorous convergence analysis of recurrence to the target distribution $\tilde{q}_{t}({\bm{x}}_{t})$.
Theorem 1.
Suppose $\tilde{q}_{t}({\bm{x}}_{t})$ has bounded support, is $\alpha$-strongly log-concave and $L$-log-smooth, and $\nabla^{2}\log\tilde{q}_{t}$ is $M$-Lipschitz. Denote ${\bm{x}}_{t}^{N_{\text{recur}}}$ as the sample after $N_{\text{recur}}$} steps of recurrence, we can bound the Wasserstein distance between the distribution of ${\bm{x}}_{t}^{N_{\text{recur}}}$ and $\tilde{q}_{t}$ as:
$W_{2}(p({\bm{x}}_{t}^{N_{\text{recur}}}),\tilde{q}_{t})=\mathcal{O}\left(\sqrt{\lambda_{t-1}-\lambda_{t}}+e^{-2\lambda_{t}}-e^{-2\lambda_{t-1}}+(1-e^{-2\lambda_{t}}+e^{-2\lambda_{t-1}})^{N_{\text{recur}}}\right)\,$,
where $\lambda_{t}=\log\frac{\alpha_{t}}{\sigma_{t}}$ is half of the log SNR.
Proof.
Recall recurrence is equivalent to the following recursion equation, with $a_{t}$, $r_{t}$ being sampler-dependent parameters in Eq. 13:
${\bm{x}}_{t}^{i+1}$ $={\bm{x}}_{t}^{i}+a_{t}r_{t}\nabla_{{\bm{x}}_{t}}\log\tilde{q}_{t}({\bm{x}}_{t}^{i})+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}$
$={\bm{x}}_{t}^{i}+a_{t}\nabla_{{\bm{x}}_{t}}\log\tilde{q}_{t}({\bm{x}}_{t}^{i})^{r_{t}}+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}\,.$
Thus, recurrence is equivalent to running unadjusted Langevin algorithm (ULA) on the tempered distribution $p^{\text{tempered}}\propto\tilde{q}_{t}^{r_{t}}$. Using Lemma 1 and Lemma 2 from Wibisono (2018), given the regularity conditions on $\tilde{q}_{t}$, we can bound the discretization error from ULA as:
$W_{2}(p^{\text{tempered}},p({\bm{x}}_{t}^{N_{\text{recur}}}))$ $=\mathcal{O}\left(a_{t}+(1-a_{t})^{N_{\text{recur}}}\right)$
$=\mathcal{O}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}+(1-\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}+\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}})^{N_{\text{recur}}}\right)$
$=\mathcal{O}\left(e^{-2\lambda_{t}}-e^{-2\lambda_{t-1}}+(1-e^{-2\lambda_{t}}+e^{-2\lambda_{t-1}})^{N_{\text{recur}}}\right)\,.$
To bound $W_{2}(p^{\text{tempered}},\tilde{q}_{t})$, we first bound the TV distance as $\text{TV}(p^{\text{tempered}},\tilde{q}_{t}$).
Denote $Z(r_{t})=\int\tilde{q}_{t}(x)^{r_{t}}dx$, and $\psi(r)=\log Z(r)$. We have:
$\text{KL}(p^{\text{tempered}},\tilde{q}_{t})$
$=\mathbb{E}_{p^{\text{tempered}}}\left[r_{t}\log\tilde{q}_{t}-\psi(r_{t})-\log\tilde{q}_{t}\right]$
$\leq\mathcal{O}\left((r_{t}-1)\mathbb{E}_{p^{\text{tempered}}}\left[\log\tilde{q}_{t}-\psi^{\prime}(r)\right]+\psi^{\prime\prime}(r)(r_{t}-1)^{2}\right)\,$,
where the last step is using Taylor expansion. Since $\psi^{\prime}(r)=\mathbb{E}_{p^{\text{tempered}}}[\log\tilde{q}_{t}]$ and $\psi^{\prime\prime}(r)=\text{Var}_{p^{\text{tempered}}}[\log\tilde{q}_{t}]$. Using the Pinsker inequality, we have $\text{TV}(p^{\text{tempered}},\tilde{q}_{t})\leq\frac{1}{2}\sqrt{\text{KL}(p^{\text{tempered}},\tilde{q}_{t})}\leq\mathcal{O}\left((r_{t}-1)\sqrt{\text{Var}_{p^{\text{tempered}}}[\log\tilde{q}_{t}]}\right)=\mathcal{O}(r_{t}-1)$, given the finiteness of $\text{Var}_{p^{\text{tempered}}}[\log\tilde{q}_{t}]$ under bounded support.
Following Proposition 7.10 in Villani (2003) for distributions with bounded support, we have:
$W_{2}(p^{\text{tempered}},\tilde{q}_{t})$
$=\mathcal{O}\left(\sqrt{\text{TV}(p^{\text{tempered}},\tilde{q}_{t})}\right)$
$=\mathcal{O}\left(\sqrt{r_{t}-1}\right)$
$=\mathcal{O}\left(\sqrt{1-\min\left(\frac{\alpha_{t}\sigma_{t-1}}{\alpha_{t-1}\sigma_{t}},\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\right)}\right)$
$=\mathcal{O}\left(\sqrt{\log\frac{\sigma_{t}}{\sigma_{t-1}}+\max\left(\log\frac{\alpha_{t-1}}{\alpha_{t}},\log\frac{\sigma_{t}}{\sigma_{t-1}}\right)}\right)$
$=\mathcal{O}\left(\sqrt{\log\frac{\alpha_{t-1}}{\alpha_{t}}+\log\frac{\sigma_{t}}{\sigma_{t-1}}}\right)$
$=\mathcal{O}\left(\sqrt{\lambda_{t-1}-\lambda_{t}}\right)\,$,
where we recall the definition of $r_{t}$ in Eq. 13.
Putting together the bound on $W_{2}(p^{\text{tempered}},p({\bm{x}}_{t}^{N_{\text{recur}}}))$ and $W_{2}(p^{\text{tempered}},\tilde{q}_{t})$, we obtain our desired bound. ∎